Let E be an event such that P(E) 0.420 and D be the event such that P(D)-0.89 Determine P(E)
Answer #1

E' is the complimentary event of E

Hence P(E) + P(E') = 1

Given

P(E) = 0.42

P(E') = 1 - P(E)

= 1 - 0.42 = 0.58

P(E') = 0.58